OpenStudy (anonymous):

what is the simplified form of 9y^5/15x^7 times 5x^3/45y^2 help!!!

OpenStudy (anonymous):

the options are a) 15x^3/y^4 b) 15x^4/y^3 c) y^4/15x^3 d) y^3/15x^4

OpenStudy (helder_edwin):

do you remember this? \[ \frac{a}{b}\frac{c}{d}=\frac{ac}{bd} \] and \[\large \frac{a^n}{a^m}=a^{n-m}=\frac{1}{a^{m-n}} \]

OpenStudy (anonymous):

not really..... can you explain it to me?

OpenStudy (helder_edwin):

OK! you have \[ \large \frac{9y^5}{15x^7}\cdot\frac{5x^3}{45y^2} \] when you multiply fractions you do it numerator times numerator and denominator times denominator so you get what?

OpenStudy (anonymous):

i thought you cant do that cause they arent like terms?

OpenStudy (helder_edwin):

yes you dont mix unlike terms but since you have multiplication, u just put them together

OpenStudy (helder_edwin):

so you have \[ \large \frac{(9y^5)(5x^3)}{(15x^7)(45y^2)} \]

OpenStudy (helder_edwin):

now \[ 9\times5=45 \] so you can cancel out 9 and 5 in the numerator with 45 in the denominator, right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

wouldnt you cancel out the 45 too?

OpenStudy (helder_edwin):

so you now have \[ \large \frac{y^5x^3}{15x^7y^2} \] can you apply the rules i gave at the beginning?

OpenStudy (anonymous):

so the answer is y^3/ 15x^4 ?

OpenStudy (anonymous):

aka D

OpenStudy (helder_edwin):

Yes!

OpenStudy (anonymous):

yaaay thank you